OpenStudy (jazzyfa30):

Solve the equation by using the Square Root Property. x^2 = 21 - 4x

hartnn (hartnn):

do you know how to complete the square ?

OpenStudy (jazzyfa30):

no

hartnn (hartnn):

x^2+4x = 21 here, the co-efficient of 'x' is 4 so, divide it by 2, and take its square, what u get ?

OpenStudy (jazzyfa30):

4/2=2 square root is 1.4

hartnn (hartnn):

we need to SQUARE the result so, 2^2 = 4 now, we add this on both sides., x^2+4x+4 = 21+4 = 25 got this ?? notice something about left side ?

OpenStudy (jazzyfa30):

u added 4????

hartnn (hartnn):

yes, i added the result (here, we got 4) to both sides now notice that left side is a perfect square! isn't it ?

OpenStudy (jazzyfa30):

yes

hartnn (hartnn):

x^2+4x+4 is square of what ?

OpenStudy (jazzyfa30):

2

hartnn (hartnn):

4 is the square of 2 x^2+4x+4 will be square of \(\large (x+2) \\ (x+2)^2 =x^2+4x+4\) so we have \((x+2)^2=25\) take square root of both sides , what u get ?

OpenStudy (jazzyfa30):

x+1.4=5

OpenStudy (jazzyfa30):

x+2=5 my b

hartnn (hartnn):

yes.. just put +/- in front of 5 x+2=5 or x+2=-5

hartnn (hartnn):

can you solve those 2 easy equations ? x+2= 5....x =.. ? x+2=-5, ....x=....?

OpenStudy (jazzyfa30):

x=3 anx =-7

hartnn (hartnn):

correct! good work :)

OpenStudy (anonymous):

JAZZY ITS BRONX

OpenStudy (jazzyfa30):

hi what happenedy u not on ur regulrs account

OpenStudy (jazzyfa30):

Solve the equation by factoring: x2 = 18 - 3x

OpenStudy (goldphenoix):

Does x2 = x^2 And does the = sign suppose to be a + sign?

OpenStudy (jazzyfa30):

sorry im still waking up x^2=18-3x

OpenStudy (goldphenoix):

I don't think you can factor that.

OpenStudy (jazzyfa30):

sorry @hartnn i fell asleep i was up all night literally i didnt geta ounce of sleep until after w finished the problem we just finished

hartnn (hartnn):

\(x^2+3x-18=0\) can you tell me 2 numbers with sum = 3 and product = -18 ?

OpenStudy (jazzyfa30):

6 and -3

hartnn (hartnn):

get some rest then! you can continue in the morning ? yes! 6 and -3 are correct :)

OpenStudy (jazzyfa30):

it is morning hear its 8:30 am

hartnn (hartnn):

so, split up, 3x = 6x-3x \(x^2+6x-3x-18=0\) factor out x from 1st 2 terms and -3 from last 2.....can you ?

OpenStudy (jazzyfa30):

(x+6)-(x-3) i think

hartnn (hartnn):

there's no - *minus* sign \((x+6)(x-3)=0\) is correct final answer! good :)

OpenStudy (jazzyfa30):

oh yeah im sorry yall

hartnn (hartnn):

and since you need to solve find x from x+6=0 and x-3=0

OpenStudy (jazzyfa30):

-6 andd 3

hartnn (hartnn):

\(\large \checkmark \)

OpenStudy (jazzyfa30):

thank you 5 more and im going to sleep

hartnn (hartnn):

ok! how about asking new question in new post? :)

OpenStudy (jazzyfa30):

i did